Going from point A to point B a driver averages 60mph. On the return trip along the same route he averages 40mph. What is the average rate for the entire trip?
Found 2 solutions by mananth, Alan3354:
Answer by mananth(12270)

A to B = 60 mph
B to A = 40 mph
..
lets be practical
while going to B
speed = 40 mph
Assume it takes 1 hour
so distance = 40 miles
..
while returning the speed = 60 mph
distance = 40
time= 40/60 = 2/3 hours
..
Average speed for the trip = total distance / total time
..
Total distance = 80 miles
total time= 1+2/3 hours
=1 2/3
=5/3 hours
..
80/(5/3)
80*3/5
=16*3
=48 mph is the average speed
..
does your teacher approve of this logic
..
m.ananth@hotmail.ca



Answer by Alan3354(30993)

It's not 50 mph. More time is spent moving at 40 mph than moving at 60 mph. It's always less than the average of the 2 speeds, unless they're equal.
-------------
d = rt
t = d/r
d = the distance, so t = 2d/r for the round trip
2t = d/60 + d/40 = 100d/2400 = d/24 (total time)
t = d/48
r = 48 mph